# Agglomerative Clustering of Simulation Output Distributions Using Regularized Wasserstein Distance

### v1

@@ 4.3 Monitoring Queue Lengths
In this experiment, we illustrate an approach that enables a decision maker to monitor the system in real time and use offline clustering to make staffing adjustments. In our setup, the system consists of 22 basic service, 9 premium service, and 8 technical operators, with other system specifications remaining the same as before. The state of the system is represented by a 4-dimensional vector consisting of the queue lengths for regular and premium customers for initial service, and the queue lengths for regular and premium customers for technical service. For a given state, we consider three KPIs: the mean utilization of all operators, the sum of the maximum waiting times of regular and premium customers in the technical queue, and the total number of customers who abandoned the queues (referred to as customer churn), all measured over a one-hour period when starting in that state. To construct a set of scenarios for our offline experiment, we first simulate the system for 5000 days, recording the states at the beginning of each hour along with the corresponding output vector after an hour of observation. We restrict our attention to those states that were observed 10 or more times, of which there were 113. Algorithm 2 groups the output distributions into three clusters, the barycenters of which are depicted in Figure 6. Across all three performance metrics, Cluster 1 performs the best, Cluster 2 performs moderately well, and Cluster 3 performs the worst.
We now have the tools to monitor and classify the states visited during a new day by considering the state’s two nearest neighbors, as measured in terms of the queue lengths. When the current state’s two nearest neighbors belong to Cluster 1, we anticipate good performance in the next hour. Conversely, having both nearest neighbors in Cluster 2 suggests high customer churn and moderate operator utilization, with minimal impact on maximum waiting times. If both nearest neighbors belong to Cluster 3, poor performance across all metrics is expected in the next hour. There will be cases where the nearest neighbors are of different kinds, i.e., we find ourselves in a transition state between clusters. When this is the case, the decision maker should closely monitor trends and be prepared to take preventive actions, such as adjusting staffing in bottleneck areas or reallocating roles among cross-trained operators in a call-center context.
Figure 6: Marginal cdfs for each KPI for each barycenter, for state online monitoring.
@@ 5 Conclusion
This paper introduces an efficient agglomerative clustering algorithm for multivariate empirical distributions, motivated by the setting of analyzing simulation output data. Clustering simulation output data by scenario can be a powerful approach for anomaly detection, pre-optimization, and classification in online monitoring. Future research directions include clustering simulation output distributions in a streaming-data setting and clustering simulation sample paths, which can provide deeper insights into dynamic system behavior.

### v2

@@ 4.3 Monitoring Queue Lengths
In this experiment, we illustrate an approach that enables a decision maker to monitor the system in real time and use offline clustering to make staffing adjustments. In our setup, the system consists of 22 basic service, 9 premium service, and 8 technical operators, with other system specifications unchanged. The state of the system is represented by a 4-dimensional vector consisting of the queue lengths for regular and premium customers for initial service, and the queue lengths for regular and premium customers for technical service. For a given state, we consider three KPIs: the mean utilization of all operators, the sum of the maximum waiting times of regular and premium customers in the technical queue, and the total number of customers who abandoned the queues (referred to as customer churn), all measured over a one-hour period when starting in that state. To construct a set of scenarios for our offline experiment, we first simulate the system for 5000 days, recording the states at the beginning of each hour along with the corresponding output vector after an hour of observation. We restrict our attention to those states that were observed 10 or more times, of which there were 113. Algorithm 2 groups the output distributions into three clusters, the barycenters of which are depicted in Figure 9. Across all three performance metrics, Cluster 1 performs the best, Cluster 2 performs moderately well, and Cluster 3 performs the worst.
We could subsequently monitor and classify states visited during a new day by considering the state’s two nearest neighbors, as measured in terms of the queue lengths. When the current state’s two nearest neighbors belong to Cluster 1, we anticipate good performance in the next hour. Conversely, having both nearest neighbors in Cluster 2 suggests high customer churn and moderate operator utilization, with minimal impact on maximum waiting times. If both nearest neighbors belong to Cluster 3, poor performance across all metrics is expected in the next hour. There will be cases where the nearest neighbors are of different kinds, i.e., we find ourselves in a transition state between clusters. In such cases, the decision maker should closely monitor trends and be prepared to take preventive actions, such as adjusting staffing in bottleneck areas or reallocating roles among cross-trained operators in a call-center context.
Figure 9: Smoothed pdfs for each KPI for each barycenter for state online monitoring.
@@
4.4 Common Random Numbers
In our final experiment, we investigate the effects of employing CRN on the clustering process. To implement CRN in the call center model, we modify the service times to be customer specific and pre-generate all customers along with their associated service times, their probability of requiring technical support, and their patience time. The same set of pre-generated customers are used to simulate a given day (replication) under different staffing configurations.
We first test whether CRN can reduce the variability of the regularized Wasserstein distance. We simulate 40 days under two staffing configurations, $(9,23,27)$ and $(7,28,14)$—once independently and once using CRN—and in each case calculate the regularized Wasserstein distance between the two empirical distributions. This process is repeated 100 times and the distributions of the regularized Wasserstein distance under both settings are plotted in Figure 11. The estimated variances of the regularized Wasserstein distance for the independent sampling and CRN cases are $3.26\times 10^{-5}$ and $2.15\times 10^{-5}$, respectively, and an $F$-test testing whether the two variances are equal yields a $p$-value of $0.0198$.
We next study whether CRN can affect how the clusterings produced by our algorithm in Section 4.1 compare with the true clustering of the scenarios’ output distributions, were those distributions known. To obtain a proxy for the true clustering, we simulate each staffing configuration 2000 times, apply our agglomerative clustering algorithm, and treat the result as the true clustering. For both the independent and CRN cases, we simulate each staffing configuration over 15 days, and conduct 100 macroreplications. To measure the similarity between two clusterings, we employ the Adjusted Rand Index (ARI) introduced by Hubert and Arabie (1985). The ARI ranges from $-1$ to $1$, where $1$ indicates the clusterings match perfectly. The distributions of the ARI using independent sampling and CRN are depicted in Figure 11. The mean ARI for clusterings with independent sampling is approximately 0.9461 with a standard deviation of 0.0736, whereas the mean ARI for clusterings using CRN is approximately 0.9845, with a lower standard deviation of 0.0194. These results provide some empirical evidence that, compared to independent sampling, using CRN can produce clusterings that are not only more consistent with the ground truth but also less variable.
Figure 11: (Left) distribution of regularized Wasserstein distance between the empirical distributions for staffing configurations $(9,23,27)$ and $(7,28,14)$ based on 100 macroreplications for CRN and independent sampling. (Right) Adjusted Rand Index (ARI) for clusterings resulting from CRN and independent sampling.
@@ 5 Conclusion
This paper introduces an efficient agglomerative clustering algorithm for multivariate empirical distributions, motivated by the setting of analyzing simulation output data. Clustering simulation output data by scenario can be a powerful approach for anomaly detection, pre-optimization, and online monitoring. The proposed algorithm demonstrates superior computational efficiency in terms of both the choice of distance metric (regularized Wasserstein) and the clustering methodology (agglomerative). Limited experiments suggested that synchronizing random numbers across scenarios, à la CRN, can both reduce the variability of the returned clustering and improve its accuracy. Also A Python package for the agglomerative clustering algorithm is under development. Future research directions include clustering simulation output distributions in a streaming-data setting and clustering the curves of state variables over time for discrete-event simulations, which can provide deeper insights into dynamic system behavior.